### v3

# Christian Ghiglino    Alastair Langtry

Abstract
This paper adapts ideas from social identity theory to set out a new framework for modelling conspicuous consumption. Notably, this approach can explain two stylised facts about conspicuous consumption that initially seem at odds with one another, and to date have required different families of models to explain each: (1) people consume more visible goods when their neighbours’ incomes rise, but (2) consume less visible goods when incomes of those with the same race in a wider geographic area rise. The first fact is typically explained by ‘Keeping up with the Joneses’ models, and the second by signalling models. Our model also explains related features of conspicuous consumption: that the rich are more sensitive to others’ incomes than the poor, and that the effect of income inequality on consumption differs qualitatively across groups. Importantly, it explains this fourth stylised fact without falling back on differences in preferences across groups, as required in other models. In addition, our model delivers new testable predictions regarding the role of network structure and income inequality for conspicuous consumption.
JEL:               D63, D85, D91
Keywords:      Social identity, Keeping Up with the Joneses, networks, centrality,
income inequality
Economists widely recognise that people’s consumption decisions respond to the consumption decisions of others. Existing literature documents two main stylised facts about these responses. First, an influential paper by Charles et al. (2009) documents that, holding their own permanent income fixed, people in the US consume less when average income of those of the same race and who live in the same state is higher. Second, work by Kuhn et al. (2011) shows that people consume more when average income of their geographic neighbours increases. This is often called a “Keeping up with The Joneses” effect. On the face of it, these two stylised facts may appear to conflict. Plus, the classes of models typically used to rationalise each of them – signalling models for the Charles et al. (2009) result and social comparisons models for the Kuhn et al. (2011) result – cannot capture both simultaneously.
We present a simple model of social identity that explains both of these stylised facts as a single phenomenon. In it, people derive status directly from consuming visible goods themselves, and from belonging to an identity that consumes lots of visible goods. In our model, identity group status depends on average consumption of visible goods by members of the identity group, and we assume there is some substitutability between status from each of these two sources. People also feel pressure to conform to some benchmark for their identity, and we assume it depends only on a person’s neighbours rather than their entire identity group. These two features are both core components of social identity theory (Costa-Font and Cowell, 2015). Our model is also able to speak to the first stylised fact without relying on the characteristics in the lowest part of the distribution – which is needed in signalling models – aligning our model and its predictions more closely with available data.
We also explain two further stylised features of consumption decisions. Third, Kaus (2013) finds that richer people are more sensitive to changes in the income of others. Our model predicts exactly this behaviour. [^2] A large part of the contribution in Kaus (2013) is to extend the analysis in Charles et al. (2009) to a lower income country – South Africa. He replicates the findings for Black South Africans (who make up the large majority of the South African population), but not for White South Africans. But the finding regarding sensitivity is novel to Kaus (2013). Fourth, Charles et al. (2009) find that higher dispersion of income within a person’s reference group leads to lower conspicuous consumption among Whites, but higher conspicuous consumption among Blacks. Our model is flexible enough to capture these contrasting behaviours. It can do so because it finds that the overall impact of a change in inequality depends on the network structure. Therefore we can explain both of these findings if Blacks and Whites have different network structures. We believe explaining the fourth stylised fact in this way is a useful contribution because it is able to capture contrasting behaviours without relying on differences in preferences across racial groups. This is something that existing signalling models can struggle with.
On the way to explaining these four stylised facts, we provide a simple characterisation of behaviour. Equilibrium consumption depends critically on a generalised version of Bonacich centrality (Katz, 1953; Bonacich, 1987) – a well-known measure of network centrality. But here, both an agent’s own centrality and the average centrality among her identity group are important. In line with existing work, centrality is increasing in the strength of network connections, in an agent’s own income and in the income of those who share her identity. Related to this, we find that stronger homophily – when people are more tightly connected to those like themselves – increases consumption. This is because being more connected to people who share her identity increases the pressure an agent feels to conform – which pushes up consumption.
We also study the effects of changing inequality. We consider a setting with a number of separate communities, and examine the inequality between different communities and within a given community separately. The impact of redistributing across different communities depends critically on the relative densities of network connections in the affected communities. This highlights the role that network connections play in mediating the impact of inequality. When considering the impact of redistribution within a community, we impose a regularity condition on the network to simplify the analysis. Subject to this condition, the impact depends on whether the redistribution is progressive or regressive. A progressive redistribution increases overall consumption, while a regressive redistribution does the opposite. This lines up with the standard intuition that poorer agents’ consumption choices are more responsive to changes in incomes.
The key novelty compared to existing models of conspicuous consumption is that we have two sources of status – own consumption and average consumption of the identity group – and that they are imperfect substitutes. That status comes from one’s identity group is a core idea in the social identity theory literature. But as in existing models of social comparisons, agents still feel pressure to keep up with their friends – creating a strategic complementarity between a person’s own consumption and that of her neighbours. [^3] One difference to existing models of social comparisons is that agents in our model feel pressure to conform to their friends – they dislike falling behind, but also dislike getting too far ahead – rather than feeling pressure to get ahead. But this is not important for the results. So when one agent’s consumption changes, whether this leads some other agent to raise or lower her consumption in response depends on how closely connected the two are in the network.
To see how this explains the two main stylised facts, suppose that incomes increase for some people. These people increase their consumption of visible goods – the traditional direct response to an increase in own income. Now consider a single person whose income did not change. The change in other people’s income creates two spillover effects. First, it increases the status benefits for the single person, by increasing the average consumption of those who share her identity. This then pushes her to consume fewer visible goods – because the status from her identity and the status from her consumption of visible goods are substitutes. Second, it increases her neighbours’ consumption, which increases the pressure to consume visible goods through the dissonance costs. This pushes her to consume more visible goods.
Which of the two forces wins out depends on how closely connected the single person is to the people whose incomes increase. This is precisely what delivers both of the stylised facts. When the people whose incomes increase are close neighbours (as in Kuhn et al. (2011)) we should expect the second force to win out – so she consumes more visible goods. In contrast, when these people are not closely connected to her (as in Charles et al. (2009)) we should expect the first force to win out – delivering the seemingly counter-intuitive result that she consumes fewer visible goods.
The rest of the paper is organised in the usual way. Section 1 briefly reviews related literature. Section 2 presents the model. Section 3 presents the results. Section 4 shows how our model explains the stylised facts. Then Section 5 examines the impact of inequality. Section 6 concludes.
1 Literature
Our paper sits within a large literature on interpersonal comparisons and conspicuous consumption. [^4] It spans disciplines, including marketing, sociology, social psychology (see Dubois and Ordabayeva (2017) and Dubois et al. (2021) for recent reviews), and economics (where it dates at least to Veblen (1899), with other early contributions from Duesenberry (1949); Festinger (1954); Frank (1985b, a) and Van de Stadt et al. (1985)). Bursztyn and Jensen (2017) provide an interesting survey of the wider topic of social pressure. Within economics, there are two broad views on why people engage in conspicuous consumption. The first assumes that people care about their income or wealth relative to others – and conspicuous consumption is used as a signal of high income/wealth. Charles et al. (2009) use this signalling approach to explain the first stylised fact – when others get richer, a given agent moves down the income ranking, and so has fewer people she needs to signal she is richer than (their model follows Glazer and Konrad (1996) very closely). However, with no explicit notion of ‘neighbours’, these models struggle to explain the second stylised fact – that people consume more when their neighbours’ incomes rise. Additionally, because agents are consuming to differentiate themselves from people poorer than themselves, everyone’s consumption decisions can be sensitive to the income of those at the very bottom of the distribution.
The second view assumes that people make social comparisons direction over consumption – so people gain status from conspicuous consumption itself. The core idea is that people like high consumption relative to others, not just in absolute terms – often called a “Keeping up with the Joneses” motive. It is perhaps the more dominant of the two views. Our paper is most closely related to a recent strand of this literature, started by Ghiglino and Goyal (2010), that uses explicit networks to allow for very rich patterns of social comparisons. More recent contributions that use explicit networks include Immorlica et al. (2017); Langtry (2022) and Bramoullé and Ghiglino (2022). [^5] Before explicit networks were introduced, most authors either assumed that people care about their consumption relative to the population average (Abel, 1990; Clark and Oswald, 1996; Ljungqvist and Uhlig, 2000) or about their rank within the population (Frank, 1985a, b, 2001; Hopkins and Kornienko, 2004). More recent contributions include Luttmer (2005); Heffetz (2011); Frank et al. (2014); Drechsel-Grau and Schmid (2014); Alvarez-Cuadrado et al. (2016); Jinkins (2016); Bertrand and Morse (2016) and De Giorgi et al. (2020). These ‘Keeping up with the Joneses’ models successfully explain the second stylised fact. But they cannot explain the first. This is because they embed a desire to ‘Keep up with the Joneses’, which makes agents’ consumption choices strategic complements – an increase in some people’s income must (weakly) increase everyone’s consumption.
What our paper adds is the ability to explain a set of so-far difficult to reconcile stylised facts in a single framework. It does so by building a new type of model – one motivated by social identity, rather than signalling or ‘Keeping up with the Joneses’ motives.
In doing so, we also relate to a strand of literature that examines the link between inequality and conspicuous consumption. By and large, formal models provide ambiguous results on the nature of the link. In Charles et al. (2009), a regressive redistribution of income (from a poorer agent to a richer one) increases total conspicuous consumption only if consumption is convex in income – but this is not a consistent property of observed consumption data (Heffetz, 2011). [^6] Other versions of the signalling model make similar predictions (Ireland, 2001; Heffetz, 2011; Moav and and Neeman, 2012; Jinkins, 2016). Further, Bilancini and Boncinelli (2012) show that the relationship between conspicuous consumption and inequality can be sensitive to the choice of whether to use ordinal or cardinal utility (also see Hopkins (2023) for similar discussion).
In line with this literature, we do not present an ironclad relationship between inequality and conspicuous consumption. But we do provide novel predictions regarding how the nature of the relationship depends on properties of the network. Notably, it allows us to explain our fourth stylised fact – that higher dispersion in income within a person’s reference group leads to lower conspicuous consumption among Whites, but higher conspicuous consumption among Blacks – without an appeal to differences in underlying preferences across groups, but instead different network structures across groups.
Finally, our paper relates to the burgeoning literature on social identity theory within economics (see Costa-Font and Cowell (2015) for a survey). The core idea in these models is that agents are part of an ‘identity’ group, and gain some status from belonging to the group, but also bear some costs from deviating from some benchmark for that group (either real differences, or in the agents’ own perception of matters). This setup was formalised in seminal contributions from Akerlof and Kranton (2000) and Shayo (2009). [^7] A number of recent papers extend or apply Shayo’s approach; see for example, Klor and Shayo (2010); Gennaioli and Tabellini (2019); Grossman and Helpman (2021); Lindqvist and Östling (2013), and Shayo (2020). These papers allow people to choose their identity group, but keep each identity’s prototype (the benchmark for ‘normal’) exogenously fixed.
Our main contribution here is to show how models of social identity can help resolve a puzzle regarding conspicuous consumption behaviour – something these models have not been used for to date. From a more technical standpoint, we also make the benchmark that agents find it costly to deviate from (often called a ‘prototype’ in the literature) endogenous. In effect, this demands that there are not persistent differences between actual behaviour in equilibrium and the benchmark for ‘normal’ behaviour. Our utility formulation closely follows Shayo (2009) but with an endogenous ‘prototype’. However, we abstract away from the choice of identity group – which is often a key focus in this literature. This is both in the interest of tractability and because it is not needed to explain the stylised facts we focus on.
2 Model
Agents and endowments.
There are $J$ agents, with typical agent $j$. Each agent is endowed with income $w_{j}>0$ and a fixed identity $\theta_{j}$, from a finite set of possible identities $\Theta$. Agents are embedded in a weighted and directed network represented by a $J\times J$ non-negative matrix $G$. We say that agent $k$ is a neighbour of $j$ if $G_{jk}>0$, and assume that $j$ is not her own neighbour (so $G_{jj}=0$ for all $j$). Let $I(\theta)=\{k:\theta_{k}=\theta\}$ be the set of agents who have identity $\theta$.
Consumption game.
Agents simultaneously choose consumption of a visible good, $x_{j}\geq 0$.
Reference point.
Each agent $j$ has a reference point, $R_{j}=\sum_{k\in I(\theta_{j})}G_{jk}x_{k}$; a weighted sum of the consumption of her neighbours who share her identity.
Preferences.
Agent $j$ gets status benefits that are increasing in her own consumption, $x_{j}$, and in the average consumption of agents with the same identity as her, $Y(\theta_{j})=\frac{1}{|I(\theta_{j})|}\sum_{k\in I(\theta_{j})}x_{k}$. And we assume that the status from each of these sources are imperfect substitutes: the marginal benefit of each is decreasing in the level the other. [^8] We also assume that agent $j$ ignores her own impact on $Y(\theta_{j})$ when choosing her consumption. This has a vanishingly small impact as $J$ grows large.
Agent $j$ also experiences dissonance costs when her consumption differs from her reference point, and these costs are convex. Note that $j$’s status benefits depend on everyone who shares her identity, but her reference point depends only on her neighbours who share her identity. There is also a convex cost of own consumption, with both absolute and marginal cost decreasing in own income. This cost function captures spending on other goods (outside of our model) in a reduced form way; implicitly assuming concave benefits from consuming the unmodelled goods.
In order to fix ideas and to show the results more clearly, we impose a particular functional form. We assume agents have the following preferences:
[…]
with $\alpha>0$, $\beta>0$, $\gamma>0$. This strong functional form assumption allows for more transparent analysis, especially regarding the role of the network.
Information, Timing, Solution Concept.
Each agent knows her own preferences, the average consumption who share her identity, and the consumption choices of her neighbours. All agents move simultaneously. So we look for Nash Equilibrium of the game.
2.1 Machinery & Assumptions
The preferences in Eq (1) put our model into the class of linear-quadratic games on a network (see Jackson and Zenou (2015); Bramoullé and Kranton (2016) for reviews). As is standard in this class of games, some variant on Bonacich centrality will play an important role in equilibrium behaviour. And to ensure an equilibrium exists we will need the largest eigenvalue of the network to be strictly less than one (Ballester et al., 2006; Bramoullé et al., 2016).
Our variant on Bonacich centrality is intuitively much the same as the standard definition, but uses only within-identity links in the network, and these links are weighted by the importance of the dissonance function, $\beta$, and by agents’ incomes, $w_{j}$. The reason that only the within-identity links play a role is because we assumed that in both the status and dissonance functions, an agent only cares about those who share her identity. Formally, define a network $H$ where $H_{jk}=\frac{\beta w_{j}}{\beta w_{j}+1}\widehat{G}_{jk}$ for all $j,k$, with $\widehat{G}_{jk}=G_{jk}\cdot\mathbf{1}\{\theta_{j}=\theta_{k}\}$ for all $j,k$ (where $\mathbf{1}\{\cdot\}$ is the usual indicator function). It will be important to bear in mind that the network $H$ is a function of $\widehat{G}$, agents’ incomes, and of $\beta$.
Definition 1 (Generalised Bonacich Centrality).
The centrality of agent $j$ is:
$C^{b}_{j}=\sum_{k}\left[\sum_{t=0}^{\infty}H^{t}\ \right]_{jk}\cdot\frac{w_{k}}{\beta w_{k}+1}$ (2)
In the interest of brevity, we will just refer to this as ‘centrality’. It will also be convenient to define the average centrality of agents with a given identity. Formally, let $\bar{C}^{b}_{\theta}=\frac{1}{|I(\theta)|}\sum_{k\in I(\theta)}C^{b}_{k}$ be the average centrality of agents with identity $\theta$.
Note that our eigenvalue condition applies to the network $H$ (rather than $G$). It ensures that the infinite sum in Equation 2 converges, and hence that the centrality measure is well-defined.
Assumption 1.
$|\lambda_{1}|<1$, where $\lambda_{1}$ is the largest eigenvector modulus of the matrix $H$.
Given the functional form we impose in Equation 1, in order to guarantee that status benefits are increasing in $Y(\theta_{j})$, we require that no agent’s centrality is ‘too far’ above the average. Formally, it amounts to the following (necessary and sufficient) restriction.
Assumption 2.
Assume that for all agents $j$:
$C^{b}_{j}<\frac{1}{\gamma}\cdot\frac{f(\bar{C}^{b}_{\theta_{j}})^{3}}{\gamma+\alpha f(\bar{C}^{b}_{\theta_{j}})},\ \text{ where }\ f(\bar{C}^{b}_{\theta_{j}})=\frac{\alpha}{4}\bar{C}^{b}_{\theta_{j}}+\frac{1}{4}\cdot(\alpha\bar{C}^{b}_{\theta_{j}})^{0.5}\cdot(\alpha\bar{C}^{b}_{\theta_{j}}+\frac{8\gamma}{\alpha})^{0.5}.$
Intuitively, each agent’s centrality must be less than some (increasing) function of the average centrality. It is clear that the higher the average centrality, the more dispersion is permitted.
We impose Assumptions 1 and 2 throughout the paper, and will not restate them for each result.
2.2 Discussion
Our utility formulation in Equation 1 closely follows Shayo (2009) – its two core components are the status benefits and dissonance costs. Shayo (2009), as in many other models, focuses on how people choose their identity. Therefore he treats the dissonance costs as depending only on the identity choice – not on actions. Given our different focus, we make the opposite choice: agents’ reference points – the ‘benchmark for normal’ that they want to conform to – is determined by people’s endogenous consumption choices. Note that in the social identity literature, the ‘benchmark for normal’ is typically called a prototype (including in Shayo (2009, 2020)), rather than a ‘reference point’. [^9] The ‘reference point’ terminology is more common in the literature on social comparisons, and stems from that literature’s more explicit link to behavioural economics.
Further, we allow the reference point to depend on an agent’s position in a network. This follows the now-common notion that people compare themselves to those they actually interact with (e.g. Immorlica et al. (2017); Bramoullé and Ghiglino (2022)). The trade-off is that we treat identity as exogenously fixed in order to keep the model tractable and ensure clean results.
Note that in line with typical social identity models, agents dislike being away from the reference point here – either above or below. This takes the view that people want to conform. It contrasts with the standard assumption in the social comparisons literature that people want to get ahead – and so dislike being below the reference point, but enjoy being above it. Nevertheless, it is possible to replace the social identity style desire to conform with a social comparisons style desire to get ahead. This change does not alter the key insights or the qualitative features of equilibrium behaviour. We show this in the Online Appendix. [^10] To remain highly tractable, it requires a linear, rather than quadratic, cost of consumption.
Further, empirical evidence in favour of each of these two views of how agents feel about being above the reference point is weak. In contrast, the idea that agents dislike being below the reference point—an idea common to both social identity and social comparisons models—finds strong empirical support.
Before showing how our model rationalises the stylised facts set out in the introduction, we first analyse the model on its own terms. We characterise equilibrium behaviour, and show how it depends critically on an agent’s own centrality, and the average centrality of her identity group. Then we examine how the primitives of the model impact centrality.
3 Results
We start by solving the model: that is providing an exact characterisation of how agents behave in equilibrium.
Proposition 1.
There exists a unique equilibrium. For all $j$, equilibrium consumption is:
$x_{j}^{*}=\frac{1}{2}\left(\alpha+\frac{\gamma}{f(\bar{C}^{b}_{\theta_{j}})}\right)C_{j}^{b}.$ (3)
Note that $f(\bar{C}^{b}_{\theta_{j}})$ is defined as in 2. Centrality is the key determinant of equilibrium consumption. As in other linear-quadratic games, higher own centrality increases an agent’s own consumption. This is because agents with high centrality are more connected to others who themselves have high consumption – giving them a high reference point. In turn this pushes them to consume more in an effort to conform to the ‘benchmark for normal’ set by the reference point.
There is also high status from belonging to a group where the other members high consumption (which is driven by them having high centrality). But status from the group is imperfectly substitutable with the status frown own consumption. So when an agent belongs to a high status group, she has a lower marginal benefit of own consumption – so she does less of it. The substitutability between the two sources of status is key. When an agent belongs to a high status group, she feels less need to create status for herself through her own consumption. The following result formalises this discussion.
Remark 1.
$x_{j}^{*}$ is strictly increasing in $C_{j}^{b}$, and strictly decreasing in $\bar{C}^{b}_{\theta_{j}}$, for all $j$.
This links equilibrium behaviour to centrality. The next natural question is to ask what drives centrality. High centrality in our model comes from a combination of strong connections to neighbours (especially to those who are themselves well connected), connections to people with high income, and having high income oneself.
Remark 2.
$C_{j}^{b}$ is increasing in $w_{\ell}$, and in $G_{\ell m}$, for all $j,\ell,m$. These results are strict if and only if $j$ and $\ell$ share an identity and there is a walk from $\ell$ to $j$.
More and stronger links raise an agent’s reference point – the benchmark she tries to conform to. This follows from the definition of the reference point. So she raises her consumption in response. In turn, this then pushes up her neighbours’ reference points, which raises their consumption, and so on. The effect ripples out through the network. Higher income makes consumption cheaper for an agent (importantly, at the margin) – so she chooses to consume more directly. This then raises her neighbours’ reference points, and so ripples out through the network in much the same way as with stronger links. In both cases, the ripple effect is contained within the affected identity group – other identity groups are unaffected.
This is because we assumed that the reference point – the ‘benchmark for normal’ that agents feel pressure to conform with – depends only on those with the same identity. It captures that idea that I only feel pressure to conform to ‘people like me’. Those with a different identity are seen as different, and so not conforming with them does not present any issues.
A consequence of this feature is that homophily has a particularly clean impact on centrality. Homophily is the tendency for people to be disproportionately connected to those similar to themselves, and is one of the most ubiquitous features of social networks (Jackson, 2008; Pin and Rogers, 2016). Higher homophily means more links between people with the same identity (at the expense of fewer cross-identity links). And because it is only the within-identity links that matter for our centrality measure, higher homophily has much the same impact as higher network density.
To show this properly, we now need a formal definition of homophily. There are many in the literature, but we use the homophily index as one of the simplest available. The homophily index is specific to an agent and simply measures the proportion of total links that are with agents in the same group. And we naturally take the identity groups as the groups of interest.
Definition 2 (Homophily index: from Currarini et al. (2009)).
For an agent $j$ in group $\theta$:
$\phi_{j}=\frac{\sum_{k\in I(\theta)}G_{jk}}{\sum_{k}G_{jk}}$ (4)
Additionally, we need to look at some particular type of changes to the network. We go for the simplest possible: swapping a link $G_{jk}$ for a link $G_{j\ell}$. [^11] It is clear that any change to the network structure that holds network density (i.e. the weighted total number of links) constant can be created from a series of link swaps. It is clear from 2 that changing a link $G_{jk}$ to a link $G_{j\ell}$ only affects the homophily index of agent $j$. So, when examining the effect of changing a link, we can conflate changes in the homophily index for agent $j$ and homophily more broadly without problems. It also follows easily from the definition that homophily falls when $j,k$ share an identity and $\ell$ has a different one, that homophily rises when $j,\ell$ share an identity and $k$ has a different one, and is left unchanged otherwise. [^12] To see this, first notice that the denominator on the right-hand side of eq. 4 never changes – this is because we are only changing links. So homophily falls if changing the link reduces the number of within-identity links. Similarly, homophily rises if changing the link increases the number of within-identity links. These other cases are that (a) $j,k$ and $\ell$ all share an identity, and (b) $j$ has a different identity to both $k$ and $\ell$. In case (b) it does not matter whether or not $k$ and $\ell$ have the same identity because the link between $k$ and $\ell$ (if it exists) is not affected.
In this setting, we can then say that changing links in a way that increases homophily will increase centralities. And similarly, changing links in a way that decreases homophily will decrease centralities.
Remark 3.
Swap a link $G_{jk}$ for a link $G_{j\ell}$.
(i) if homophily rises, then $C_{m}^{b}$ rises for all $m\in I(\theta_{j})$,
(ii) if homophily falls, then $C_{m}^{b}$ falls for all $m\in I(\theta_{j})$.
This result is driven by the fact that only within-identity links that matter for equilibrium behaviour (recall that it is the matrix $\widehat{G}$ rather than the matrix $G$ that appears in the definition of centrality). This is because it is only these links that drive the reference point. In our model, agents do not care per se about the decisions of those from different identity groups. So changing a link in the network $G$ in a way that increases homophily adds a links in the network $\widehat{G}$ – which is the one that enters into the definition of centrality. [^13] Note that this observation also means that swapping a between-identity link for a different between-identity link has no effect. Swapping two different within-identity links is more complicated. There, both adding the link $G_{j\ell}$ and removing $G_{jk}$ have an impact on the network $\widehat{G}$. However, high centrality is driven by being linked to people who themselves have high centrality. So $j$’s centrality rises if swapping the link connects her to someone who has higher centrality. Conversely, if the swap connects her to someone with lower centrality, her own centrality falls. This follows straightforwardly from Langtry (2022, Proposition 3).
The effect of discrimination.
In reality, the status of different identity groups will depend on a number of other (exogenous) factors in addition to the visible good consumption of the group’s members that this paper focuses on. [^14] It would be straightforward to add it to the model. Simply define $Y(\theta_{j})=\frac{1}{|I(\theta_{j})|}\sum_{k\in I(\theta_{j})}x_{k}+E(\theta_{j})$, where $E(\theta_{j})>0$ is some exogenous parameter capturing status from other (exogenous factors). This exogenous status component may be linked to discrimination or marginalisation in a society. Groups that are discriminated against or marginalised would have a low exogenous status component, while favoured groups would have high exogenous status.
Importantly, this will affect their consumption choices. All else equal, agents belonging to the marginalised groups will be pushed towards greater visible good consumption—and away from non-visible good consumption. [^15] We do not model other goods explicitly, but the convex cost function captures the opportunity cost of foregoing consumption of non-visible goods in stylised manner. This is because they use visible good consumption to compensate for their otherwise low status. This follows from one of the key features of our model – that agents get status both from their identity group and from their own consumption, and, crucially, that these two sources are imperfect substitutes.
The broader impacts of this will depend on the properties of visible goods compared to non-visible ones. While what constitutes a visible (status generating) good will depend on context, the standard view of such goods is as luxury consumption goods (e.g. Charles et al. (2009)). Non-visible goods cover everything else, but importantly include durable goods and investments in capital. So discrimination and marginalisation pushed agents away from spending on non-visible goods, and by extension may push them away from investments in capital.
This suggests a new potential harm of discrimination: it pushes agents towards greater visible good consumption, at the loss of investment in non-visible capital goods. To our knowledge, this is a novel prediction about how discrimination and marginalisation affect behaviour. Formal analysis of welfare effects is beyond the scope of this paper – we would need to model non-visible goods explicitly and examine how agents benefit from them.
4 The effects of changes in income
In the introduction, we set out a series of stylised facts about how changes in incomes feed through into conspicuous consumption decisions. And we noted that existing models struggle to rationalise them all as a coherent whole. The main aim of this section is therefore to show how our model fits all of these facts simultaneously.
In the previous section we found that a change in incomes affects both agents’ centralities and the average centralities of their identity groups – and that these have opposing effects on equilibrium behaviour. While well-known intuitions regarding Bonacich centrality style measures suggest that the impact of a change to incomes (or other primitives) have a larger impact on agents closer in the network to the change. So the key idea is that for agents located ‘close’ to the change, the effect on own centrality will dominate the effect on average centrality. And that the reverse applies for agents located ‘far from’ the change. But it is difficult to state precise results for general network structures.
To show this clearly we will therefore focus on a stark benchmark setting where agents are split up into disconnected ‘communities’. This will allow us to separate each channel cleanly. And it gives us a stylised way of thinking about ‘near’ and ‘far’ in the network. A pair of agents who share a community are ‘near’ to one another in the network, and a pair in different communities are ‘far apart’.
Definition 3 (Communities network).
The network $\widehat{G}$ consists of $N$ strongly connected components, [^16] A walk in $\widehat{G}$ from $j$ to $k$ is a sequence of agents $j,j^{\prime},...,k^{\prime},k$ such that $\widehat{G}_{j^{\prime}k^{\prime}}>0$ for all $j^{\prime},k^{\prime}$ in the sequence. A strongly connected component is a set of agents such that for all agents $j,k$ in the set, there exists a walk from $j$ to $k$. with typical component $n$, and where $N$ is large. Call each component a community.
Changes in one community will only affect centralities within that one community, and will not spill over into other communities. This means a change in one community affects other communities only through the impact on average centrality. Further, assuming that there are many communities guarantees that a change in community $n$ has a larger impact on the centrality of the agents in that community than it does on the average centrality of all agents. In summary, following a change in community $n$, the ‘own centrality’ channel dominates within community $n$, and the ‘average centrality’ channel dominates in all other communities.
Before stating our result, it is helpful to quickly recap the stylised facts that relate to changes in income. This will help make the link between them and the behaviour predicted by our model easier to see.
  1. 1.
    In the US, people consume less visible goods when the average income of others of the same race in their state is higher (Charles et al., 2009).
  2. 2.
    In the Netherlands, people consume more when the average income of their geographic neighbours increases (Kuhn et al., 2011).
  3. 3.
    In South Africa, people’s consumption of visible goods is more sensitive to changes in income of their reference group when they are richer (Kaus, 2013).
Note that we defer the fourth stylised fact to later in the section. This is because it relates to changes in inequality, and is less clear-cut – and so needs dedicated analysis. With these facts in mind, we can now state our result.
Proposition 2.
Consider a communities network. Suppose income increases for some agents with identity $\theta$ in community $n$.
(i)   Consumption decreases for all agents with identity $\theta$ in communities other than $n$,
(ii)  Consumption increases for all agents with identity $\theta$ in community $n$.
(iii) For all agents $j$ in communities other than $n$, the marginal impact on $j$’s consumption is increasing in $j$’s own income.
The mapping from 2 into the stylised facts is fairly direct. An increase in incomes in other communities corresponds to higher average income of others of the same race in the same state, as in Charles et al. (2009). [^17] We use the term ‘income’ in our model for convenience. Where appropriate, this should be taken to mean the ‘permanent income’ examined by Charles et al. (2009) in the empirical analysis. An increase in incomes within an agent’s own community corresponds to raising her neighbours’ incomes, as in Kuhn et al. (2011). Note that 2 applies only to agents who share an identity, but Kuhn et al. (2011) does not consider identity in their analysis. Nevertheless, strong homophily and sorting along geographic lines are well-documented phenomena within economics, sociology McPherson et al. (2001); Schelling (2006) and other disciplines. So it is very likely that at least some of an agent’s geographic neighbours do in fact share her identity. The third stylised fact is even more direct: an ‘increasing marginal impact’ is simply a formalisation of the ‘greater sensitivity’ found by Kaus (2013).
From a more technical standpoint, 2 follows from 1 when focusing attention on the communities network. In that case, an increase in the incomes of some agents with identity $\theta$ in community $n$ does two things. First, it increases the centralities of agents with identity $\theta$ in community $n$. Second, this then increases the average centrality of identity $\theta$. Importantly, it does not affect the centrality of identity $\theta$ agents in other communities. So consumption falls in the communities other than $n$ – because average centrality rose, with no impact on own centrality. And consumption rises in community $n$ – because the effect on own centrality dominates the effect on average centrality. The sensitivity result is also immediate: higher own income increases own centrality, and this acts multiplicatively, making consumption more sensitive to other changes.
More importantly, 2 illuminates the mechanism driving this behaviour. An increase in incomes for some agents with identity $\theta$ in community $n$ raises their consumption directly. This then raises the reference points of their neighbours, raising the benchmark these neighbours are trying to conform to – and hence raising their consumption. This effect ripples out through the whole community (see discussion following 2). This then raises the average consumption of agents with identity $\theta$. So agents with that identity gain more status from belong to that identity group. So those in other communities then feel less need to seek status through their own consumption of visible goods.
But for agents in community $n$, the changes are concentrated ‘close’ to them in the network. So the pressure to conform to a higher benchmark is strong relative to the gains from their whole identity group having higher status. It is this pressure from a higher benchmark that pushes up consumption – and this is the force that dominates inside community $n$.
While Propositions 1 and 2 are clear on how the third stylised fact – the greater sensitivity of agents with higher incomes – is captured mechanically, the headline intuition is relates to the way income affects the costs of consumption. In particular, agents with higher income have a lower marginal cost of consumption, and this marginal cost rises more slowly. When the marginal status benefits of consumption rise, as happens when $Y(\theta_{j})$ decreases, it should not come as a surprise that agents increase their consumption in response. But by how much they adjust their consumption depends on how quickly the marginal cost rises as consumption rises. It is precisely higher income agents whose marginal costs rise more slowly, and so who make greater adjustments.
5 The effects of changes in inequality
Having examined changes in the level of income, we now turn to changes in the dispersion of income; that is, changes in inequality. When focusing on a setting with distinct communities (as we did in Section 4 and continue to do so here), there are two dimensions to income inequality: inequality between different communities, and inequality within a given community. In the interest of presenting clean results, we will examine these dimensions one by one. Each dimension yields complementary insights.
Looking at between-community inequality (where we abstract away from within-community inequality) highlights the role of network density. There, communities with denser networks are more sensitive to changes in their income, So redistributing incomes from one community to another increases total consumption when the receiving community has a denser network. And this then pushes communities not party to the redistribution to cut their consumption. This is driven by the substitutability between own consumption and identity group status discussed earlier.
Looking at within-community inequality (where we impose some additional structure on the network in the interest of tractability) highlights how inequality affects consumption. An increase in inequality reduces total consumption, but raises consumption in the communities where income did not change. Again, this is driven by the substitutability between the sources of status.
5.1 Between community inequality
In order to focus on the between-community dimension of inequality, we will assume here that all agents within a given community have the same income. Formally, that $w_{i}=w_{n}$ for all $i\in n$. With this in place, we can now define the ‘density’ of a community. This will be a key metric when studying the effects of changing between-community inequality.
Definition 4 (Community Density).
The density of community $n$ is:
$D_{n}=\sum_{j,k\in n}\left[\sum_{t=0}^{\infty}\widehat{G}^{t}\right]_{jk}$
Readers familiar with Bonacich centrality will notice that our definition of ‘density’ is the sum of centralities of all agents in the community (using the ‘standard’ definition of Bonacich centrality that only involves the network, rather than our variant that also includes income). It is not network density as conventionally defined. [^18] See Jackson (2008, S.2.1.8) for the conventional definition. Nevertheless, it captures an important aspect of how dense connections in a community are on aggregate.
We focus on a particularly simple form of redistribution – a transfer from one community to another, keeping average income in society constant. It reduces income inequality if the community receiving the transfer is poorer than the one giving the transfer, and similarly increases income inequality if the receiving community is richer. We will not consider the ambiguous case where the transfer changes which community is richer. It is clear that much richer changes to the income distribution can be built from a series of these changes.
Proposition 3 (Inequality).
Consider two communities, $n$ and $n^{\prime}$, both with identity $\theta$. And consider a transfer of income from $n$ to $n^{\prime}$ such that the ranking of the two communities’ incomes does not change.
  • (i)
    Consumption falls in community $n$ and rises in community $n^{\prime}$,
  • (ii)
    Total consumption rises if and only if $D_{n^{\prime}}>D_{n}$,
  • (iii)
    Consumption on all other islands rises if and only if $D_{n^{\prime}}<D_{n}$.
The first part of this result is unsurprising. When everyone in a community has higher income, they consume more. And they consume less when they have lower income. But interestingly, the impact of this transfer on overall consumption does not depend on whether it increases or decreases inequality. Rather, it depends on the relative densities of the networks within the two communities that are party to the transfer. If the community receiving the transfer has a denser network than the community giving the transfer, then overall consumption rises. This is because centrality is higher in the receiving community. Therefore consumption there is more sensitive to a change in income than in the giving community. So the increase in consumption in the receiving community more than offsets the fall in the giving community.
This raises total consumption across the two communities – in turn raising the status of the identity group. As before, other (unaffected) communities respond to an increase in the status of their identity groups by cutting their own consumption – due to the substitutability between the two sources of status. But note that this cut in consumption by the other (unaffected) communities cannot outweigh the initial rise in consumption across the two affected communities.
5.2 Within community inequality
In order to focus on the within-community dimension of inequality, we impose a regularity condition on the network structure. Specifically, we assume that each agent has the same number of neighbours (in the weighted sense) with those of each different income level. For example, if there are two different levels of income, $w_{1}$ and $w_{2}$, then each agent has $L$ neighbours with income $w_{1}$, and $L$ neighbours with income $w_{2}$. [^19] Note that if every agent in the network has a unique income level, then this assumption requires that the network is complete. But we believe that there being a small number of different levels of income is a reasonable first-order approximation of reality. Note that this is stronger than assuming the network is regular (which would only require that each agent has $L$ links in total, regardless of the neighbours’ incomes). Formally,
Assumption 3.
Let community $n$ have $Z$ different levels of income, $\{w_{1},...,w_{Z}\}$. Then $\sum_{k}\widehat{G}_{jk}\cdot\mathbf{1}\{w_{k}=w_{m}\}=L$ for all $z$ and for all $j\in n$.
Intuitively, 3 ensures that how strongly $j$ is connected to others is unrelated to how much income those other people have. Colloquially, she is, on average, connected to “Mr and Mrs Average”. This retains the network effects, but “smooths them out” so that they apply equally throughout the network.
Again, we focus on a the simplest possible form of redistribution. Here, that is transfer of income from agents with income $w_{z}$ to agents with income $w_{z^{\prime}}$. This is the same as our transfer in the previous subsection, except that now the transfer happens between sets of agents inside the same community – rather than between two different communities. It reduces income inequality if the agents receiving the transfer are poorer than the agents giving the transfer (and similarly increasing inequality if the receiving agents are richer). As before, we will not consider the ambiguous case where the transfer changes which set of agents is richer.
Proposition 4.
Consider a community $n$. And consider a transfer of income from agents with (initial) income $w_{z}$ to agents with (initial) income $w_{z^{\prime}}$, such that the ranking of the two groups’ incomes does not change.
  • (i)
    Consumption falls for agents with income $w_{z}$ and rises for agents with initial income $w_{z^{\prime}}$,
  • (ii)
    Total consumption rises if and only if the transfer reduces inequality (i.e. $w_{z^{\prime}}<w_{z}$),
  • (iii)
    Consumption of all other agents falls if and only if the transfer reduces inequality (i.e. $w_{z^{\prime}}<w_{z}$).
With the regularity condition on the network, we can show that consumption is increasing and concave in own income. So the first part of this result is standard. But this concavity also drives the second part. It means that a given change in income has a larger impact on lower income agents. So a progressive (inequality reducing) transfer raises overall consumption – it increases consumption of the poor by more than it reduces the consumption of the rich. This increase in consumption has a spillover effect for everyone else in the same identity group: higher average consumption increases the status from belonging to the identity group. This increase in status induces agents who were not party to the transfer to cut their own consumption in response – due to the substitutability between status from own consumption and status from the identity.
The key insight is that transfers that affect inequality can have important impacts on agents whose incomes do not change. This is because agents derive status from consumption by others who share their identity. And this ‘status identity’ is substitutable with status derived from their own consumption – so lower status identity leads them to compensate with more of their own consumption. The result apes existing results in the conspicuous consumption literature, but our mechanism is novel.
5.3 Explaining an empirical feature
Beyond their results concerning how income levels affect consumption of visible goods, Charles et al. (2009) also document a puzzling empirical regularity concerning the effect of income inequality. They find that higher inequality among Whites in a given US state leads to lower consumption of visible goods by Whites, holding own income constant. But they find the opposite for Blacks: higher inequality among Blacks in a given US state leads to higher consumption of visible goods by Blacks, holding own income constant.
We can rationalise these qualitatively different responses to changing inequality across racial groups as perhaps being due to a non-monotonic relationship between community density and community income. More precisely, suppose that density (a la 4) follows a U-shaped pattern: it is first decreasing and then increasing in income. Note that density relates to how people set the ‘benchmark for normal’ that they want to conform to, and hence how strongly people make social comparisons. This non-monotonicity essentially assumes that low and high income agents/communities feel the strongest pressures to conform.
Under this assumption, an increase in inequality among only high income communities will involve shifting income from lower density communities to higher density ones. 3(iii) tells us that consumption in the communities whose income did not change will then rise. Conversely, an increase in inequality among only low income communities will involve shifting income from higher density communities to lower density ones. In this case, 3(iii) tells us that consumption in the communities whose income did not change will then fall. Figure 1 provides an illustration.
[…]
Figure 1: Gray line shows a possible relationship between income and density in different communities. Blue crosses (left panel) show possible realisations for Black communities. Green pluses (right panel) show possible realisations for White communities. Red dots (both panels) show a change due to a regressive transfer between two communities.
If such a non-monotonic relationship between density and income exists, then we should reasonably expect Blacks to be disproportionately below the peak of the curve, and Whites to be disproportionately above it. This is simply due to the well-documented fact that Blacks have, on average, lower incomes than White (Charles et al., 2009). So an increase in inequality among White would raise consumption of visible goods by Whites (among those whose income did not change), and the reverse would happen for Blacks. This matches exactly the finding by Charles et al. (2009).
Given that (often strong) homophily along many different dimensions is an extremely robust finding in the empirical literature on social networks (Currarini et al. (2009),McPherson et al. (2001),Currarini and Mengel (2016), Jackson et al. (2022)), we contend it is plausible that agents make stronger comparisons with those who are more similar to themselves in terms of location and income. If so, then the non-monotonic relationship that we posit in Figure 1 could be consistent with a pair of known stylised facts.
First, Bailey et al. (2020) show that higher income agents have more geographically dispersed social networks – proportionately less of their friends are nearby. Second, Chetty et al. (2022) show that there is substantial homophily based on socioeconomic status (which is closely tied to income) – people are disproportionately friends with those who have similar income to themselves. Importantly, they find that this homophily is substantially stronger at the upper end of the income distribution. Within the confines of our model, this would push richer agents to have stronger social comparisons, and so higher centrality.
Within the confines of our model, the first of these stylised facts would push richer agents to have weaker social comparisons, and so lower centrality. The second would do the opposite – push richer to have stronger social comparisons, and so higher centrality. Importantly, the increasing geographic dispersion (first fact) appears to hold for the whole income distribution, while the increasing homophily along income lines (second fact) appears to only kick in towards the top of the income distribution. Taken together, these could generate the non-monotonicity posited in Figure 1.
Therefore, our model can rationalise this puzzling empirical finding without relying on differences across racial groups, beyond the well-documented differences in incomes. One drawback of our approach is that it relies on the density of agents’ social networks, which is likely to be challenging to measure empirically. Nevertheless, we view it as a useful contribution because existing signalling models struggle to explain this empirical finding without appealing to racial differences in preferences.
6 Conclusion
A key feature of our model is that we make the prototype and reference point – what is viewed as ‘normal’ for an identity group – endogenous. This departs from the standard assumption in social identity theory, and so avoid situations where what is considered ‘normal’ for an identity group permanently differs from what is actually normal for an identity group. However, this comes at the cost of having to hold identity fixed. While in the context of what we are exploring, we believe this trade-off to be a very rewarding one, it is nevertheless a limitation of the model. Models that can allow both an endogenous/reference point and endogenously chosen identity, while retaining tractability, would represent a significant contribution to this literature. We have shown the importance of an endogenous prototype, and Shayo (2009) (among others) has shown the importance of endogenous identity choice. Achieving both within a tractable model is a material gap in existing literature.
Appendix A Proofs
Proposition 1
First, we find the best response function, holding $Y(\theta_{j})$ constant (the first derivative is linear and decreasing in $x_{j}$, and is positive when $x_{j}=0$, so a unique solution exists). Note that we have substituted in $R_{j}=\sum_{k\in I(\theta_{j})}G_{jk}x_{k}=\sum_{k}\widehat{G}_{jk}x_{k}$.
$x_{j}$ $=\frac{1}{2}\left(\frac{\beta w_{j}+1}{w_{j}}\right)^{-1}[\alpha+\gamma Y(\theta_{j})^{-1}+2\beta\widehat{G}_{jk}x_{k}]$
$\implies x_{j}-\frac{\beta w_{j}}{\beta w_{j}+1}\widehat{G}_{jk}x_{k}$ $=\frac{1}{2}\frac{w_{j}}{\beta w_{j}+1}[\alpha+\gamma Y(\theta_{j})^{-1}]$
Then recall that we defined the matrix $H$ such that $H_{jk}=\frac{\beta w_{j}}{\beta w_{j}+1}\widehat{G}_{jk}$. We now substitute this in to the equation and switch to matrix notation. In an abuse of notation use the same notation for a vector of constants as for the equivalent scalar, and we let $\mathbf{W}$ denote the vector where the $j^{th}$ entry is $\frac{w_{j}}{\beta w_{j}+1}$. And recall that $\odot$ is the Hadamard, or element-wise, product. Further, because we have assumed that $|\lambda_{1}|<1$ (1), [^20] Where $\lambda_{1}$ is the largest eigenvalue modulus (i.e. the largest in magnitude) of the matrix $H$., so $(I-H)$ is invertible [Ballester et al., 2006, Theorem 1]. Therefore:
$\mathbf{x}-H\mathbf{x}$ $=\frac{1}{2}\mathbf{W}\odot[\alpha+\gamma Y(\mathbf{\theta})^{-1}]$
$\implies\quad\qquad\mathbf{x}$ $=(I-H)^{-1}\cdot\frac{1}{2}\mathbf{W}\odot[\alpha+\gamma Y(\mathbf{\theta})^{-1}].$
Reverting to scalar notation:
$x_{j}=\sum_{k}[(I-H)^{-1}]_{jk}\cdot\frac{1}{2}\frac{w_{k}}{\beta w_{k}+1}[\alpha+\gamma Y(\theta_{k})^{-1}].$
Now notice that, by construction, $\widehat{G}$ is a block matrix with off-diagonal blocks being all zero. This is because an agent $j$ can only have a link to agent $k$ if $j$ and $k$ share the same identity (i.e. if $\theta_{j}=\theta_{k}$). Therefore $H$ is also a block matrix. As a result, $H_{jk}$ will be zero whenever $j$ and $k$ have different identities (i.e. whenever $\theta_{j}\neq\theta_{k}$). An important consequence of this is that $\theta_{k}=\theta_{j}$ whenever $[(I-H)^{-1}]_{jk}>0$. This means we can treat the term $Y(\theta_{k})^{-1}$ as a constant an pull it out of the sum. So
$x_{j}=\frac{1}{2}[\alpha+\gamma Y(\theta_{j})^{-1}]\cdot\sum_{k}[(I-H)^{-1}]_{jk}\cdot\frac{w_{k}}{\beta w_{k}+1},$
and then using 1, and the fact that $(I-H)^{-1}=\sum_{t=0}^{\infty}H^{t}$ (this is the Neumann series representation of a matrix inverse), gives us:
$x_{j}=\frac{1}{2}[\alpha+\gamma Y(\theta_{j})^{-1}]C_{j}^{b}.$
Recall that $Y(\theta_{j})=\frac{1}{|I(\theta_{j})|}\sum_{k\in I(\theta_{j})}x_{k}$, and so is endogenous. Hence, by substituting our equation for consumption into this definition:
$Y(\theta_{j})$ $=\frac{1}{|I(\theta_{j})|}\sum_{k\in I(\theta_{j})}\frac{1}{2}[\alpha+\gamma Y(\theta_{j})^{-1}]C_{k}^{b}$
$Y(\theta_{j})$ $=\frac{1}{2}[\alpha+\gamma Y(\theta_{j})^{-1}]\cdot\frac{1}{|I(\theta_{j})|}\sum_{k\in I(\theta_{j})}C_{k}^{b}.$
The term after the dot is, by definition, $\bar{C}^{b}_{\theta_{j}}$ (the average Bonacich centrality of agents with identity $\theta_{j}$, see definition immediately after eq. 2). Using this definition and rearranging terms yields:
$0\displaystyle 2Y(\theta_{j})^{2}-\alpha\bar{C}^{b}_{\theta_{j}}Y(\theta_{j})-\gamma\bar{C}^{b}_{\theta_{j}}=$ 0
Then apply the usual quadratic formula and ignore the negative root (because we rule out $x_{j}<0$ by assumption and hence cannot have $Y(\theta_{j})<0$).
$Y(\theta_{j})=\frac{1}{4}\alpha\bar{C}^{b}_{\theta_{j}}+\frac{1}{4}(\alpha\bar{C}^{b}_{\theta_{j}})^{0.5}\cdot(\alpha\bar{C}^{b}_{\theta_{j}}+\frac{8\gamma}{\alpha})^{0.5}$
Finally, substitute in this equilibrium value of $Y(\theta_{j})$ into our equation for $x_{j}^{*}$.
$x_{j}^{*}=\frac{1}{2}\left(\alpha+\frac{\gamma}{\frac{1}{4}\alpha\bar{C}^{b}_{\theta_{j}}+\frac{1}{4}(\alpha\bar{C}^{b}_{\theta_{j}})^{0.5}\cdot(\alpha\bar{C}^{b}_{\theta_{j}}+\frac{8\gamma}{\alpha})^{0.5}}\right)C_{j}^{b}$
Remark 1
Follows immediately from 1. ∎
Remark 2
For convenience, recall that $C_{j}^{b}=\sum_{k}\sum_{t=0}^{\infty}[H^{t}]_{jk}\cdot\frac{w_{k}}{\beta w_{k}+1}$. First, notice that $\frac{w_{k}}{\beta w_{k}+1}$ is increasing in $w_{\ell}$ when $k=\ell$, and is otherwise unaffected. This weakly increases $C_{j}^{b}$.
Second, a well-known property of the matrix inverse $(I-H)^{-1}=\sum_{t=0}^{\infty}[H^{t}]$ is that $\sum_{t=0}^{\infty}[H^{t}]_{jk}$ is increasing in $H_{\ell m}$, and strictly so if and only if there exists a walk (in $H$) from $j$ to $\ell$ [^21] To see this, express $\sum_{t=0}^{\infty}[H^{t}]=H^{0}+H^{1}+H^{2}+H^{3}+...$ It is immediate that all elements of $\sum_{t=0}^{\infty}[H^{t}]$ are weakly increasing in $H_{ji}$ for all $j,i$. Additionally, notice that $[H^{2}]_{j\ell}=\sum_{k}H_{jk}\cdot H_{ki}$. Therefore, $[H^{2}]_{j\ell}$ is positive and strictly increasing in $H_{k\ell}$ if $H_{jk}>0$ (i.e. there is a walk of length one from $j$ to $k$). So $\sum_{t=0}^{\infty}[H^{t}]_{jk}$ is strictly increasing in all of $k$’s links when $j$ has a walk of length 1 to $k$. Similarly, $[H^{t+1}]_{j\ell}=\sum_{k}[H^{t}]_{jk}\cdot H_{k\ell}$, and $[H^{t+1}]_{j\ell}$ is positive and strictly increasing in $H_{k\ell}$ if $[H^{t}]_{jk}>0$ (i.e. there is a walk of length $t$ from $j$ to $k$). So if there is a walk of some length (some $t>0$) from $j$ to $k$, then $\sum_{t=0}^{\infty}[H^{t+1}]_{j\ell}$ is strictly increasing in $H_{\ell m}$ for all $m$. By definition, $H_{\ell m}=\frac{\beta w_{\ell}}{\beta w_{\ell}+1}\cdot G_{\ell m}\cdot\mathbf{1}\{\theta_{\ell}=\theta_{m}\}$. So $H_{\ell m}$ is increasing in $w_{\ell}$ and in $G_{\ell m}$. Therefore $\sum_{t=0}^{\infty}[H^{t}]_{jk}$ is increasing in $w_{\ell}$ and $G_{\ell m}$, and strictly so if and only if there exists a walk (in $H$) from $\ell$ to $j$.
Note that $H_{\ell m}$ is strictly increasing in $G_{\ell m}$ whenever $\theta_{\ell}=\theta_{m}$. But for $H_{\ell m}$ to be strictly increasing in $w_{\ell}$, we need $\widehat{G}_{\ell m}>0$. This need not be true for any given $m$. But if there exists a walk from $\ell$ to $j$, then it must be the case that $\widehat{G}_{\ell m}>0$ for some $m$ (at a minimum, the link used to create the walk to $j$). So, if there exists a walk (in $H$) from $\ell$ to $j$, an increase in $w_{\ell}$ must strictly increase $H_{\ell m}$ for some $m$, and so must strictly increase $\sum_{t}[H^{t}]_{jk}$.
Finally, for there to be a walk in $H$ from $\ell$ to $j$, we must have $\theta_{\ell}=\theta_{j}$, because $H$ cannot contain any cross-identity links (by construction, it contains only within-identity links). ∎
Remark 3
Links that appear in $G$ but not in $\widehat{G}$ have no impact on consumption. Part (i) for the link swap to increase homophily, it must be the case that $\ell$ has the same identity as $j$, but $k$ does not (i.e. $\ell\in I(\theta_{j})$ and $k\notin I(\theta_{j})$). This means the link $G_{j\ell}$ appears in $\widehat{G}$, but the link $G_{jk}$ does not. So removing the link $G_{jk}$ has no impact, and this becomes equivalent to just adding $G_{j\ell}$. With this observation, the result follows from Remark 2. Part (ii) inverse of part (i). ∎
Proposition 2
First note that by 3, each community $n$ is a separate component of the network. So, for agents $j\in n$, their centrality, $C_{j}^{b}$, depends only on $\widehat{G}_{n}$ (the subgraph of $\widehat{G}$ containing only agents $j^{\prime}\in n$), not on the full network $\widehat{G}$. Next, consider an increase in $w_{j}$ for some $j\in n$ [^22] Note that because we are working with the network $\widehat{G}$, all agents in the same component must have the same identity. This is because, by design, $\widehat{G}$ only contains links between pairs of agents with the same identity. Obs. 1: it follows from 2 (and the fact that by 3, there is a walk from $j$ to $j^{\prime}$ for all $j,j^{\prime}\in n$) that $w_{j}\uparrow$ for some $j\in n$ with identity $\theta$ $\implies C_{j^{\prime}}^{b}\uparrow$ for all $j^{\prime}\in n$. Obs. 2: it is also clear that $C_{k}^{b}$ is unchanged for all $k\notin n$. Obs. 3: the fact that $C_{j^{\prime}}^{b}\uparrow$ for some $j^{\prime}\in n$ with identity $\theta$ implies that $\bar{C}^{b}_{\theta}$ increases.
Part (i) follows from Obs. 2 and Obs. 3, coupled with 1. Part (ii) follows from Obs. 1 and Obs. 3, coupled with 1 and the fact that with many different communities with identity $\theta$, the impact of the increase in $C_{j^{\prime}}^{b}$ must outweigh the impact of the increase in $\bar{C}^{b}_{\theta}$. Part (iii) The marginal effect of others’ income on an agent $j$’s consumption is $\frac{dx_{j}^{*}}{dw_{k}}$. When $j,k$ are in different communities, then $w_{k}$ does not affect $C_{j}^{b}$ (2). Using this, plus the equation for $x_{j}^{*}$ from 1 yields:
$\frac{dx_{j}^{*}}{dw_{k}}=\frac{1}{2}C_{j}^{b}\cdot\frac{d}{dw_{k}}\left(\frac{\gamma}{\frac{1}{4}\alpha\bar{C}^{b}_{\theta_{j}}+\frac{1}{4}(\alpha\bar{C}^{b}_{\theta_{j}})^{0.5}\cdot(\alpha\bar{C}^{b}_{\theta_{j}}+\frac{8\gamma}{\alpha})^{0.5}}\right).$
We know from 2 that $C_{j}^{b}$ is strictly increasing in $w_{j}$. And because there are many communities (by assumption, in 3), $w_{j}$ has a negligible impact on $\bar{C}^{b}_{\theta_{j}}$. Therefore $\frac{dx_{j}^{*}}{dw_{k}}$ is increasing in $w_{j}$. ∎
Lemma A.1.
Suppose the network is a communities network (a la 3), all communities are the same size and all agents in a given community, $P_{n}$, have the same income, $w_{n}$.
Then $C_{j}^{b}=w_{n}\cdot\sum_{k}(I-\widehat{G})^{-1}_{jk}$ for all $j\in P_{n}$.
Proof.
Start with the definition of Bonacich centrality (1). First, notice that under 3, each community is a disconnected component in the network $\widehat{G}$. Therefore, we can treat each community separately. So we can replace $\widehat{G}$ with $\widehat{G}_{n}$, which is the graph for community $n$ only. Second, recall that all agents in community $n$ have the same income, $w_{n}$.
$C^{b}_{j}=\sum_{k}\left[\sum_{t=0}^{\infty}\left(\frac{\beta w_{n}}{\beta w_{n}+1}\widehat{G}_{n}\right)^{t}\ \right]_{jk}\cdot\frac{w_{n}}{\beta w_{n}+1}.$
Now the $\frac{w_{n}}{\beta w_{n}+1}$ term at the end (far right) can be pulled out the front, and the same term that appears within the infinite sum can be separated from the $\widehat{G}_{n}$ term, and can also be pulled out of the summation over $k$.
$C^{b}_{j}=\frac{w_{n}}{\beta w_{n}+1}\cdot\sum_{t=0}^{\infty}\left(\frac{\beta w_{n}}{\beta w_{n}+1}\right)^{t}\cdot\sum_{k}\left[\sum_{t=0}^{\infty}\widehat{G}^{t}\ \right]_{jk}$
Then apply the standard closed form solution for the infinite sum of a geometric series (which always exists because $\frac{w_{n}}{\beta w_{n}+1}<1$ by construction), and notice that the result (partially) cancels with the $\frac{w_{n}}{\beta w_{n}+1}$ term. This yields: [^23] Where the final inequality is by the Neumann series representation of a matrix inverse.
$C^{b}_{j}=w_{n}\cdot\sum_{k}\left[\sum_{t=0}^{\infty}\widehat{G}^{t}\ \right]_{jk}\ =\ w_{n}\sum_{k}(I-\widehat{G})^{-1}_{jk}.$
Proposition 3
First, it follows from A.1 that centrality rises in community $n^{\prime}$ and falls in community $n$. Next, we can express average Bonacich centrality as the sum over agents in each community, then summed over communities (rather than the sum over agents directly), finally divided by $J$ (the number of agents in society). That is:
$\bar{C}^{b}_{\theta}=\frac{1}{J}\sum_{n}\sum_{j\in P_{n}}w_{n}\sum_{k}(I-\widehat{G}_{n})^{-1}_{jk}.$
Then, we use the fact that income is the same for all agents in a given community $n$, and substitute in for 4. This yields:
$\bar{C}^{b}_{\theta}=\frac{1}{J}\sum_{n}w_{n}D_{n}$
Observation. It is then clear that a transfer from community $n$ to $n^{\prime}$: (1) decreases the centrality of agents in community $n$, (2) increases the centrality of agents in community $n^{\prime}$, and (3) increases average Bonacich centrality if and only if $D_{n^{\prime}}>D_{n}$.
(i) For the communities $n$ and $n^{\prime}$, the effect of own centrality dominates the effect on average centrality (because there are many communities). Centrality, and hence consumption, rises [resp. falls] for community $n^{\prime}$ [resp. $n$].
(iii) For the other communities, own centrality does not change. Only average centrality matters. The result then follows from the observation above, and the fact that consumption is decreasing in average centrality (see 1).
(ii) It is clear that the change can only affect agents with identity $\theta$. For clarity, throughout the rest of this proof we will assume that any sum over either agents or communities is restricted to those with identity $\theta$ and will drop subscripts $\theta$. As a reminder, $i,j,k,\ell$ denote agents, and $m,n$ denote communities.
Total consumption of identity $\theta$ agents, which we denote $X$, is:
$X=\sum_{i}x_{i}^{*}$ $=\sum_{m}\sum_{j\in m^{\prime}}\frac{1}{2}(\alpha+\gamma f(\bar{C}^{b})^{-1})C_{j}^{b}$ (A.1)
$=\frac{1}{2}(\alpha+\gamma f(\bar{C}^{b})^{-1})\sum_{m}w_{m}D_{m}$ (A.2)
This uses 4: $D_{m}=\sum_{j\in m}C_{j}^{b}$. Now, W.L.O.G. suppose that the transfer of income from community $n$ to $n^{\prime}$ raises income by $\epsilon>0$ for all $j\in n^{\prime}$, and reduces income by $\epsilon$ for all $j\in n$ [^24] The fact that all communities are of equal size ensures that the size of the increase in income for each $j\in n^{\prime}$ is the same as the decrease in income for each $j\in n$ So the result amounts to characterising the partial derivative of $X$ w.r.t $\epsilon$.
$\frac{dX}{d\epsilon}$ $=\frac{1}{2}(\alpha+\gamma f(\bar{C}^{b})^{-1})(D_{n^{\prime}}-D_{n})+\frac{1}{2}\gamma(-1)f^{\prime}(\bar{C}^{b})(f(\bar{C}^{b}))^{-2}\cdot\frac{d\bar{C}^{b}}{d\epsilon}\cdot\sum_{m}w_{m}D_{m}$ (A.3)
Now, it follows straightforwardly from $\bar{C}^{b}=\frac{1}{J}\sum_{m}w_{m}D_{m}$ (shown earlier in this proof) that $\frac{d\bar{C}^{b}}{d\epsilon}=\frac{1}{J}(D_{n^{\prime}}-D_{n})$. Substituting this in, and pulling out common factors yields:
$\frac{dX}{d\epsilon}$ $=(D_{n^{\prime}}-D_{n})\cdot\left[\alpha+\gamma\frac{1}{f(\bar{C}^{b})}\left(1-\frac{f^{\prime}(\bar{C}^{b})}{f(\bar{C}^{b})}\frac{1}{J}\sum_{m}w_{m}D_{m}\right)\right]$ (A.4)
Using $\bar{C}^{b}=\frac{1}{J}\sum_{m}w_{m}D_{m}$ again yields:
$\frac{dX}{d\epsilon}$ $=(D_{n^{\prime}}-D_{n})\cdot\left[\alpha+\gamma\frac{1}{f(\bar{C}^{b})}\left(1-\frac{f^{\prime}(\bar{C}^{b})}{f(\bar{C}^{b})}\bar{C}^{b}\right)\right]$ (A.5)
Some straightforward (but tedious) algebraic manipulations then show that $\frac{f^{\prime}(\bar{C}^{b})}{f(\bar{C}^{b})}\bar{C}^{b}<1$. We prove this as And since $\alpha>0$, $\gamma>0$ (by assumption), and $f(\bar{C}^{b})>0$ (recall its definition in 2), the term in square brackets is strictly positive.
Therefore it is clear that $\frac{dX}{d\epsilon}>0$ if and only if $(D_{n^{\prime}}-D_{n})>0$. ∎
Lemma A.2.
Under 3: $C_{j}^{b}=\beta\Gamma_{j}/(1-LZ\overline{\Gamma})$, where $\Gamma_{k}\equiv\frac{w_{k}}{\beta w_{k}+1}$, and $\overline{\Gamma}=\frac{1}{|n|}\sum_{j\in n}\Gamma_{j}$.
Proof. Reminder of notation. $j,k,\ell$ denote agents. $z$ denotes an income level, $Z$ is the number of different incomes levels, and $\mathcal{Z}$ denotes the set of agents with income level $z$. The result is for community $n$, so all variables here should properly have subscripts $n$, and any sum over agents should properly be over agents in community $n$. But we omit these subscripts for cleanliness.
We first prove a claim that will be used later. Claim: 3 implies that $\sum_{k}\Gamma_{k}\widehat{G}_{jk}=ZL\overline{\Gamma}$. Proof of claim: we have, $\sum_{k}\Gamma_{k}\widehat{G}_{jk}=\sum_{z}\sum_{k\in\mathcal{Z}}\Gamma_{k}\widehat{G}_{jk}=\sum_{z}\Gamma_{z}\sum_{k\in\mathcal{Z}}\widehat{G}_{jk}=L\sum_{z}\Gamma_{z}$. The first equality is trivial – it is just summing over agents of a given income level, then over different income levels, rather than summing over agents directly. The second equality follows from the fact that $\Gamma_{k}=\Gamma_{z}$ whenever $w_{k}=w_{z}$, and by definition all agents $k\in\mathcal{Z}$ have $w_{k}=w_{z}$. The third equality follows from 3 – by assumption every agent $j$ has exactly $L$ links (in the weighted sense) to agents with income $w_{z}$ (i.e. agents $k\in\mathcal{Z}$).
Then, because the number of agents with each income level is the same (W.L.O.G. denote it $q$), we have $\sum_{j\in n}\Gamma_{j}=q\sum_{z}\Gamma_{z}$. And there are therefore $q\cdot Z$ agents in the community (which is just $|n|$ by definition). So $L\sum_{z}\Gamma_{z}=L\frac{1}{q}\sum_{j\in n}\Gamma_{j}=LZ\frac{1}{q\cdot Z}\sum_{j\in n}\Gamma_{j}=LZ\overline{\Gamma}$. This proves the claim.
Characterising centrality.
We begin by restating 1, with $\Gamma_{k}\equiv\frac{w_{k}}{\beta w_{k}+1}$:
$\frac{1}{\beta}C^{b}_{j}=\sum_{k}\left[\sum_{t=0}^{\infty}\left(\Gamma\widehat{G}\right)^{t}\ \right]_{jk}\cdot\Gamma_{k}.$
We have moved the $\beta$ term to the left-hand side simply to avoid carrying it around in the calculations. We now interchange the order of summation, then expand out the infinite sum over $t$ and characterise each term. We deal with the first two terms ($t=0$ and $t=1$) individually, and then all further terms by induction.
$\mathbf{t=0}$ term: clearly $\sum_{k}[(\Gamma\widehat{G})^{0}]_{jk}\cdot\Gamma_{k}=\sum_{k}\mathbf{1}\{j=k\}\cdot\Gamma_{k}=\Gamma_{j}$. $\mathbf{t=1}$ term: $\sum_{k}[(\Gamma\widehat{G})^{1}]_{jk}\cdot\Gamma_{k}=\sum_{k}(\Gamma\widehat{G})_{jk}\cdot\Gamma_{k}=\sum_{k}\Gamma_{j}\widehat{G}_{jk}\Gamma_{k}=\Gamma_{j}\sum_{k}\Gamma_{k}\widehat{G}_{jk}$. Applying the claim above yields: $=\Gamma_{j}\cdot LZ\bar{\Gamma}$.
Induction. For $\mathbf{t=2}$:
$\sum_{k}[(\Gamma\widehat{G})^{2}]_{jk}\Gamma_{k}$ $=\sum_{k}\sum_{\ell}(\Gamma\widehat{G})_{j\ell}(\Gamma\widehat{G})_{\ell k}\Gamma_{k}\ =\sum_{k}\sum_{\ell}\Gamma_{j}\widehat{G}_{j\ell}\Gamma_{\ell}\widehat{G}_{\ell k}\Gamma_{k}\ =\Gamma_{j}\sum_{\ell}\Gamma_{\ell}\widehat{G}_{j\ell}\sum_{k}\widehat{G}_{\ell k}\Gamma_{k}$
By the claim above, $\sum_{k}\widehat{G}_{\ell k}\Gamma_{k}=LZ\overline{\Gamma}$, which is a constant (so it can be pulled out of the sum over $\ell$). And similarly, $\sum_{\ell}\Gamma_{\ell}\widehat{G}_{j\ell}=LZ\overline{\Gamma}$. The (double) application of the claim yields
$\sum_{k}[(\Gamma\widehat{G})^{2}]_{jk}\Gamma_{k}=\Gamma_{j}\cdot(LZ\overline{\Gamma})^{2}$
For $\mathbf{t=\tau>2}$: assume that $\sum_{k}[(\Gamma\widehat{G})^{\tau}]_{jk}\Gamma_{k}=\Gamma_{j}\cdot(LZ\overline{\Gamma})^{\tau}$. Finally, show for $\mathbf{t=\tau+1}$. By definition $\sum_{k}(\Gamma\widehat{G})^{\tau+1}_{jk}\Gamma_{k}=\sum_{j}[(\Gamma\widehat{G})\cdot(\Gamma\widehat{G})^{\tau}]_{jk}\Gamma_{k}$. For clarity of exposition, let $(\Gamma\widehat{G})^{\tau}_{jk}=\mathcal{A}_{jk}$. Then
$\sum_{k}(\Gamma\widehat{G})^{\tau+1}_{jk}\Gamma_{k}=\sum_{k}\sum_{\ell}\Gamma_{j}\widehat{G}_{j\ell}\mathcal{A}_{\ell k}\Gamma_{k}\ =\Gamma_{j}\sum_{\ell}\widehat{G}_{j\ell}\sum_{k}\mathcal{A}_{\ell k}\Gamma_{k}$
Using the assumption for $t=\tau$; $\sum_{k}\mathcal{A}_{\ell k}\Gamma_{k}=\Gamma_{\ell}(LZ\overline{\Gamma})^{\tau}$. Applying this yields: $\sum_{k}(\Gamma\widehat{G})^{\tau+1}_{jk}=\Gamma_{j}\cdot(LZ\overline{\Gamma})^{\tau}\sum_{\ell}\widehat{G}_{j\ell}\Gamma_{\ell}$. Applying the claim prove above once more yields $\sum_{k}(\Gamma\widehat{G})^{\tau+1}_{jk}\Gamma_{k}=\Gamma_{j}\cdot\overline{\Gamma}^{\tau+1}$. By induction, we have calculated each term in the infinite sum.
Having calculated each term in the infinite sum, we can now express it more simply: $\frac{1}{\beta}C_{j}^{b}=\Gamma_{j}(1+(LZ\overline{\Gamma})+(LZ\overline{\Gamma})^{2}+(LZ\overline{\Gamma})^{3}+...)$ Therefore, finally:
$C_{j}^{b}=\frac{\beta\ \Gamma_{j}}{1-LZ\overline{\Gamma}}.$
Proposition 4
Consider a transfer of income such that: (i) $w_{z}\uparrow$, $w_{z^{\prime}}\downarrow$, (iii) $w_{z}+w_{z^{\prime}}$ unchanged. Now notice that $\Gamma_{j}$ is increasing and concave in $w_{j}$ (by the definition of $\Gamma_{j}$). Step 1. So the transfer causes: (i) $\Gamma_{z}\uparrow$, (ii) $\Gamma_{z^{\prime}}\downarrow$, and (iii) $(\Gamma_{z}+\Gamma_{z^{\prime}})\uparrow$ if and only if $w_{z}<w_{z^{\prime}}$ [and hence $\bar{\Gamma}\uparrow$ if and only if $w_{z}<w_{z^{\prime}}$, because the transfer does not affect $\Gamma_{z^{\prime\prime}}$ for any $z^{\prime\prime}\neq z,z^{\prime}$].
Step 2. In turn these imply: (i) $C^{b}_{z}\uparrow$, (ii) $C^{b}_{z^{\prime}}\downarrow$, and (iii) $\bar{C}^{b}_{\theta}\uparrow$ if and only if $w_{z}<w_{z^{\prime}}$. Note that for (i), $C_{z}^{b}$ is increasing in both $\Gamma_{z}$ and $\bar{\Gamma}$, but these may move in different directions. However, there are many communities by assumption, so the the transfer must have a much larger impact on $\Gamma_{z}$ than on $\bar{\Gamma}$, so the effect of $\Gamma_{z}$ dominates. The same argument applies to (ii).
Step 3. Finally, recall the definition of $x_{j}^{*}$ from 1. Our findings in Step 2 in turn imply that: (i) $x_{z}^{*}\uparrow$, (ii) $x_{z^{\prime}}^{*}\downarrow$, and (iii) $x_{z^{\prime\prime}}^{*}\uparrow$ if and only if $w_{z}<w_{z^{\prime}}$. Again, point (i) [resp. point (ii)] rely on the assumption that there are many communities to ensure that the effect of the change in $C^{b}_{z}$ [resp. $C^{b}_{z^{\prime}}$] dominates the effect of the change in $C^{b}_{z^{\prime}}$.
A.1 Additional Proofs
Here, we formalise the claims we made about the necessity and sufficiency of Assumptions 1 and 2.
Lemma A.3.
(i) Our measure of Centrality (1) exists and is well defined if and only if 1 holds. (ii) In equilibrium, $\alpha x_{j}+Y(\theta_{j})+\frac{\gamma x_{j}}{Y(\theta_{j})}$ is increasing in $Y(\theta_{j})$ if and only if 2 holds.
Proof.
[…]
$\frac{dS}{dY}$ $=1-\gamma C_{j}^{b}Y^{-2}(\alpha+\gamma Y^{-1})$
Some further rearranging shows that this is greater than zero if and only if
$C_{j}^{b}$ $<\frac{1}{\gamma}\frac{Y^{3}}{(\gamma+\alpha Y)}$
Noticing that $Y^{*}(\theta_{j})=\frac{\alpha}{4}\bar{C}^{b}_{\theta_{j}}+\frac{1}{4}\cdot(\alpha\bar{C}^{b}_{\theta_{j}})^{0.5}\cdot(\alpha\bar{C}^{b}_{\theta_{j}}+\frac{8}{\alpha})^{0.5}$ completes the result (this is shown in the proof to 1). ∎
Lemma A.4.
$\frac{f^{\prime}(\bar{C}^{b})}{f(\bar{C}^{b})}\bar{C}^{b}<1$
Proof.
First, recall that $f(\bar{C}^{b})=\frac{\alpha}{4}\bar{C}^{b}+\frac{1}{4}\cdot(\alpha\bar{C}^{b})^{0.5}\cdot(\alpha\bar{C}^{b}+\frac{8\gamma}{\alpha})^{0.5}$
For Online Publication
Appendix B Keeping up with the Joneses: an alternative to Dissonance
The model in Section 2 assumes that agents feel some pressure to conform to their respective reference points, which is a benchmark for what constitutes ‘normal’. They experience dissonance costs from being either above or below their reference point. This approach is motivated by the Social Identity Theory literature, which views deviation from a reference point as in some sense costly or unpleasant, irrespective of whether this is a deviation above or both.
However, a large part of the economics literature on social comparisons makes an alternative assumption: that agents experience costs from below their reference point, but receive benefits from being above it. This is often interpreted as pressure to ‘Keep up with the Joneses’ (or more accurately, to get ahead them). The benchmark assumption is that these benefits from ‘Keeping up with the Joneses’ (with negative benefits for consumption below the reference point) are concave in own consumption. [^26] Although some papers examine loss averse preferences, where benefits are concave when own consumption is above the reference point, but convex below. We show that our main results are robust to this alternative assumption.
So we replace the quadratic dissonance cost function in the equation capturing agents’ preferences with a square root benefit of consuming above the reference point. In order to keep the model clean, we also make two simplifying assumption. One, we assume that income is the same for everyone, so $w_{j}=w$ for all $j$. Two, we assume the cost of consumption is linear. Note that with this functional form the social comparisons component of utility is only well-defined when $x_{j}\geq Y_{j}$. This is true in equilibrium, and so does not pose a problem. So an agent’s preferences are now:
[…]
We must have $(\alpha-\frac{1}{w})<0$ for the problem to be interesting – otherwise actions clearly explode to infinity. Analogously to the main text, we will also assume that $|\lambda_{1}(\widehat{G}|<1$, where $\lambda_{1}(\widehat{G}$ is the largest eigenvalue modulus of the matrix $\widehat{G}$. And define Bonacich centrality: $\mathbb{C}_{j}^{b}=\sum_{k}(I-\widehat{G})^{-1}_{jk}$. Note that here we are using the ‘standard’ definition on Bonacich centrality. This is due to the simplifying assumption that incomes are equal for all agents.
With this setup and machinery, we can again characterise equilibrium behaviour.
Proposition B.1.
There exists a unique equilibrium. For all $j$ equilibrium consumption is:
$x_{j}^{*}=\frac{1}{4}\left(\frac{\beta}{z}\right)^{2}\mathbb{C}_{j}^{b}\left(1-\frac{2\gamma z}{2\gamma z+\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}+\frac{1}{4}(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}})^{0.5}\cdot(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}}+16\gamma z)^{0.5}}\right)^{-2}.$ (B.2)
The four key features of equilibrium behaviour are the same as in the main model. First, consumption is linearly increasing in own centrality. Second, it is decreasing in the average centrality of those who share the same identity as $j$. Third, own centrality acts multiplicatively on the term containing average centrality. Fourth, the definition of centrality depends only on the within-identity links (i.e. it depends on $\widehat{G}$, rather than on $G$). Note that the definition of centrality here is simpler because we have assumed all agents have the same income.
The presence of these four key features means that the other results from Section 3 will hold here. The exception is the first part of 2, because incomes no longer feature in the definition of centrality.
This variant on the model can therefore also explain the stylised facts. This is suggestive that it is the key novelty of the model – the multiple sources of status and the substitutability between them – that is driving the results, and not the peculiarities of the dissonance function.
B.1 Proof
The proof here follows 1 fairly closely. To make the derivations less cumbersome, define $z\equiv\frac{1}{w}-\alpha$. Now consider the marginal utility of consumption:
$\frac{du_{j}}{dx_{j}}$ $=\gamma Y(\theta_{j})^{-1}+\frac{1}{2}\beta(x_{j}-Y_{j})^{-0.5}-z$
Notice that $\frac{1}{2}\beta(x_{j}-Y_{j})^{-0.5}\geq 0$ whenever it is well-defined. So we must have $\gamma Y(\theta_{j})^{-1}-z\leq 0$ for a best response to exist. This implies that we must have $Y(\theta_{j})\geq\gamma z^{-1}$. Given this, we can now consider $j$’s best response.
$(x_{j}-Y_{j})^{-0.5}$ $=\frac{2}{\beta}z-\gamma Y(\theta_{j})^{-1}$
$\implies x_{j}-\sum_{k}\widehat{G}_{jk}x_{k}$ $=\frac{\beta^{2}}{4}\left(z-\gamma Y(\theta_{j})^{-1}\right)^{-2}$
Since we must have $Y(\theta_{j})\geq\gamma z^{-1}$ (argued above), the Right-Hand Side of the equation above is weakly positive. This implies that $x_{j}^{*}>Y_{j}$, ensures that preferences are well-defined in equilibrium. We now switch to vector notation (in an abuse of notation we use the same notation for a vector of constants as for the equivalent scalar).
$\mathbf{x}-\widehat{G}\mathbf{x}$ $=\frac{\beta^{2}}{4}\left(z-\gamma Y(\theta_{j})^{-1}\right)^{-2},$
$\mathbf{x}$ $=(I-\widehat{G})^{-1}\left[\frac{\beta^{2}}{4}\left(z-\gamma Y(\theta_{j})^{-1}\right)^{-2}\right].$
We have assumed $|\lambda_{1}|<1$, so $(I-\widehat{G})$ is invertible (Ballester et al. [2006, Theorem 1]). Revering to scalar notation:
$x_{j}=\sum_{k}(I-\widehat{G})^{-1}_{jk}\frac{\beta^{2}}{4}\left(z-\gamma Y(\theta_{j})^{-1}\right)^{-2}.$
A comment: by construction, $\widehat{G}$ is a block matrix with off-diagonal blocks being all zero. This is because an agent $j$ can only have a link to agent $k$ if $j$ and $k$ share the same identity (i.e. if $\theta_{j}=\theta_{k}$). An important consequence of this is that $\theta_{k}=\theta_{j}$ whenever $[(I-\widehat{G})^{-1}]_{jk}>0$. This means we can treat the term $Y(\theta_{k})^{-1}$ as a constant an pull it out of the sum. And we can use the (standard) definition of Bonacich centrality. So
$x_{j}^{*}=\frac{\beta^{2}}{4}\left(z-\gamma Y(\theta_{j})^{-1}\right)^{-2}\mathbb{C}_{j}^{b}$
This would give equilibrium actions for a fixed value of the prototype. But $Y(\theta_{j})$ is endogenous. Recall that by definition $Y(\theta_{j})=\frac{1}{|I(\theta_{j})|}\sum_{k\in I(\theta_{j})}x_{k}$. Hence, by substituting our equation for consumption into this definition:
$Y(\theta_{j})$ $=\frac{1}{|I(\theta_{j})|}\sum_{k\in I(\theta_{j})}\frac{\beta^{2}}{4}\left(z-\gamma Y(\theta_{j})^{-1}\right)^{-2}\mathbb{C}_{j}^{b}$
$\implies Y(\theta_{j})$ $=\frac{\beta^{2}}{4}\left(z-\gamma Y(\theta_{j})^{-1}\right)^{-2}\bar{\mathbb{C}}^{b}_{\theta_{j}},$
$\implies\frac{4}{\beta^{2}}\frac{Y(\theta_{j})}{\bar{\mathbb{C}}^{b}_{\theta_{j}}}$ $=\left(z-\gamma Y(\theta_{j})^{-1}\right)^{-2}$
using the fact that ¯$\bar{\mathbb{C}}^{b}_{\theta_{j}}$ is the average of $\mathbb{C}_{k}^{b}$ for agents $k\in I(\theta_{j})$. The LHS of the equation above is clearly monotonically increasing in $Y(\theta_{j})$. And in the region $Y(\theta_{j})\geq\gamma z^{-1}$, the RHS is monotonically decreasing in $Y(\theta_{j})$. [^27] This is less obvious, but follows from the fact that (1) $z-\gamma Y(\theta_{j})^{-1}\geq 0$, (2) this term is increasing in $Y(\theta_{j})$, and (3) this term is raised to a negative exponent. So if a solution that satisfies $Y(\theta_{j})\geq\gamma z^{-1}$ exists, it will be unique. Now rearrange terms:
$\frac{\beta^{2}}{4}\cdot\frac{\bar{\mathbb{C}}^{b}_{\theta_{j}}}{Y(\theta_{j})}$ $=z^{2}-2\gamma z\frac{1}{Y(\theta_{j})}+\frac{\gamma^{2}}{Y(\theta_{j})^{2}}$
$\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}Y(\theta_{j})$ $=z^{2}Y(\theta_{j})^{2}-2\gamma zY(\theta_{j})+\gamma^{2}$
$\implies 0$ $=z^{2}Y(\theta_{j})^{2}-\left(2\gamma z+\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}\right)Y(\theta_{j})+\gamma^{2}$
Then solve for $Y(\theta_{j})$ using the standard quadratic formula.
$Y(\theta_{j})^{=}\frac{1}{2z^{2}}\left[2\gamma z+\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}\pm\sqrt{\left(2\gamma z+\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}\right)^{2}-4z^{2}\gamma^{2}}\right]$
$Y(\theta_{j})^{=}\frac{1}{2z^{2}}\left[2\gamma z+\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}\pm\frac{1}{4}(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}})^{0.5}\cdot(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}}+16\gamma z)^{0.5}\right]$
Aside: now we show that the smaller of the two roots is not permitted, because it is less that $\gamma z^{-1}$ (which we have shown above it not possible). Suppose not:
$\frac{1}{2z^{2}}\left[2\gamma z+\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}-\frac{1}{4}(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}})^{0.5}\cdot(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}}+16\gamma z)^{0.5}\right]$ $>\gamma z^{-1}$
$2\gamma z+\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}-\frac{1}{4}(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}})^{0.5}\cdot(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}}+16\gamma z)^{0.5}$ $>2\gamma z$
$\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}$ $>\frac{1}{4}(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}})^{0.5}\cdot(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}}+16\gamma z)^{0.5}$
$\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}$ $>\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}+\psi,$
for some $\psi>0$. Contradiction. So we ignore the smaller root. Therefore:
$Y(\theta_{j})=\frac{1}{2z^{2}}\left[2\gamma z+\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}+\frac{1}{4}(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}})^{0.5}\cdot(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}}+16\gamma z)^{0.5}\right]$
Finally, we substitute this express for $Y(\theta_{j})$ back into our expression for $x_{j}^{*}$. This yields:
$x_{j}^{*}=\frac{1}{4}\left(\frac{\beta}{z}\right)^{2}\mathbb{C}_{j}^{b}\left(1-\frac{2\gamma z}{2\gamma z+\frac{\beta^{2}}{4}\bar{\mathbb{C}}^{b}_{\theta_{j}}+\frac{1}{4}(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}})^{0.5}\cdot(\beta^{2}\bar{\mathbb{C}}^{b}_{\theta_{j}}+16\gamma z)^{0.5}}\right)^{-2}.$